# Residual Connections Harm Abstract Feature Learning in Masked Autoencoders

### v2

@@ 3.2 Implementation Strategy
Skip connections for Autoencoders. Since our method progressively decays the residual connections over network depth, it encourages the most abstract features to be learned by the final layer. Hence, when using our strategy on an Autoencoder architecture, we only apply residual decay to the encoder layers. However, learning a highly abstract bottleneck is detrimental to the training objectives that aim for pixel-wise reconstruction, as they necessitate the preservation of information at the bottleneck across all feature levels. To address this, we incorporate standard skip connections between the encoder and decoder, enabling the encoder to directly pass information from shallow layers to the decoder while learning increasingly abstract representations in the deeper encoder layers.
@@
Stabilizing training with residual zero initialization. The model exhibits rapid feature norm growth at the beginning of training for $\alpha_{\rm min}\leq 0.7$. We suspect that the model learns to amplify the output feature norm of $f_{{\bm{\theta}}_{l}}({\bm{x}})$ to counteract the significant decay applied to the residual connection. This growth leads to training instability and negatively impacts training convergence. To address this issue, we follow the implementation of previous works [19] and initialize the weights of the final output layer in each $f_{{\bm{\theta}}_{l}}$ to zero instead of using the original Xavier uniform initialization [56]. This approach significantly enhances training stability by limiting the rate of feature norm growth and enables us to explore training with even lower values of $\alpha_{\rm min}$.
4 Large-scale Experiments
@@
Masked Autoencoders. We implement our method within a Masked Autoencoder (MAE) [1] by replacing the residual connections in the encoder’s MLP and attention blocks with decayed identity shortcuts. The MAE operates by accepting images with a random subset of pixels masked out and learning to recover the discarded pixels. As Section 3.2 describes, we add skip connections between the encoder and decoder to facilitate learning abstract representations at the bottleneck. Since the original MAE has twice the number of encoder layers as decoder layers, we inject output from every other encoder layer into the corresponding decoder layer. To match spatial dimensions, injected encoder features are combined with learnable masked tokens before channel-wise concatenation. The implementation details for the training and evaluation are shown in Section A.
Experimental setup. To evaluate the learned features, we first pre-train the MAE on the ImageNet-1K dataset [3] and then conduct linear probing on a learned class token from pre-trained features over the test set using a single linear layer. We also conduct ablations on ImageNet-100. For fair comparison, our baseline methods are (1) the original MAE as well as (2) our framework with $\alpha_{\rm min}=1$, which is equivalent to an MAE with skip connections. For MAE hyperparameters, including the learning rate schedule, total training epochs, and mask ratio, we follow the best settings found in the original paper. Please see the appendix in supplementary materials for detailed experimental setups.
4.1 ImageNet-1K
Method Backbone Objective Augmentation Accuracy
MoCo-v2[57] ResNet-50 Contrastive Loss Full 71.1
MoCo-v3[58] VIT-B/16 Contrastive Loss Full 76.7
SimCLR-v2[59] ResNet-50 Contrastive Loss Full 77.7
DINO[15] VIT-B/16 Contrastive Loss Full 78.2
iBOT[40] VIT-B/16 Contrastive Loss Full+Mask 79.5
Data2Vec[60] VIT-B/16 Feature Loss Full+Mask 68.0
CAE[39] VIT-B/16 Recon. + Feature Loss Affine+Mask 70.4
I-JEPA[61] VIT-B/16 Feature Loss Affine+Mask 72.9
MAE[1] VIT-B/16 Recon. Loss Affine+Mask 67.8
Ours ($\alpha_{\rm min}$ = 0.6) VIT-B/16 Recon. Loss Affine+Mask 72.7
@@ 4.1 ImageNet-1K
Table 1: Accuracy of linear classifier based on pre-trained learned representations on the ImageNet-1K dataset. We list the training objective and data augmentation process of each method. Training objectives include contrastive losses, feature losses, and image reconstruction. Affine data augmentation represents random flip, random crop, and resize. Full augmentation combines affine augmentation and color distortion. With our simple modification, we substantially improve the MAE baseline by encouraging the model to learn an increased abstract features over depth. We achieve competitive performance compared to I-JEPA which uses an explicit feature loss.
@@
We evaluate our learned representation using the ImageNet-1k dataset [3] and report the results in Table 1 by training a linear classifier on top of frozen backbone features. In the table, we show the linear probing performance of various self-supervised methods, which we categorize by their architecture, objectives, and data augmentation processes. In the top half of the table, we present methods that employ a contrastive loss. Although these methods produce the best probing accuracies, their success depends on a carefully designed data augmentation process, which may need to be tuned for each different data distribution. In the bottom half, we show several methods, including ours, which do not rely on contrastive objectives. With the exception of Data2Vec, these methods only use a standard random affine data transformation (with random masking), which need not be distribution-reliant. Among these methods, MAE only uses a pixel-wise loss, I-JEPA and CAE use a latent feature alignment loss, and CAE uses both. Our method simply extends MAE by constructing an implicit feature bottleneck and shows significant improvements over the MAE baselines (72.3% vs. 67.3%), outperforming Data2Vec and CAE and giving a probing accuracy competitive with I-JEPA, without needing to use explicit feature alignment.
(a) Residual Connection
(b) $\alpha_{\rm min}=0.6$ (ours).
@@
Figure 2: Comparison of t-SNE visualization for models trained with (a) standard Residual Connection and (b) our method with $\alpha_{\rm min}=0.6$. We use the models trained in Section 4.2 on ImageNet-100 to produce features for test samples. The visualization of t-SNE embeddings is performed on a subset comprising 10 classes selected at random. Representative images for each category along with their plot color are given on the right. We observe that the embeddings of some categories for the model trained with standard residual connections (a) have collapsed together while our method (b) forms well structured clusters.
@@
Embedding visualization. We qualitatively evaluate the feature learning in Figure 2 by visualizing the t-SNE embeddings of the learned features for a random subset of test images. While the Residual connection (baseline) embeddings in Figure 2(a) produce some separation, they struggle to differentiate some harder categories, such as tubs (brown), stairs (yellow), and church buildings (purple). In contrast, the embeddings produced from our features in Figure 2(b) display much clearer separation and are able to separate almost all categories distinctly.
4.2 Ablation
@@
Table 2: Linear Probing Accuracy on ImageNet-100 for our method varying $\alpha_{\rm min}$ and architecture size. We conduct ablation studies and demonstrate that linear probing performance for both architectures increases as $\alpha_{\rm min}$ decreases until around 0.5-0.6. While the larger ViT-B/16 architecture achieves the highest accuracy of 83.6, it is noteworthy that the smaller ViT-S/16, when utilizing our method, outperforms the baseline setting (standard Residual Connection at $\alpha_{\rm min}=1$) of ViT-B/16 MAE.
@@
Architecture size. In Table 2, we also train both the ViT-B/16 and the smaller ViT-S/16 backbone using varying $\alpha_{\rm min}$. Our framework is especially effective on the smaller architecture, increasing linear probing performance by 9.4%, compared to the baseline setting (standard Residual Connection at $\alpha_{\rm min}=1$). This is consistent with the observation that larger models generalize better [28] than smaller models. Our method is able to significantly improve the smaller ViT-S/16 and slightly close the gap between the differently-sized architectures.
Skip connections. Another critical design choice in our network is to include skip connections that are not in the original MAE. As discussed in Section 3.2, if the MAE does not use skip connections that give the network a shortcut to pass low-level detail information from the encoder to the decoder, the bottleneck layer must preserve enough information to reconstruct the input image accurately. This is opposed to the design of the shortcut decay, which prompts the bottleneck to learn abstract representations. These contrary effects significantly degrade the representation learned by the model, leading to a 22.1% drop in the linear probing score, as we report in Table LABEL:tab:skip_connections.
Different decay schema. In Section 3.1, we define our shortcut decay as a decreasing weight over every residual connection in the encoder. In Table LABEL:tab:decay_schema, we apply our framework to different network components within the encoder. Specifically, we test restricting shortcut decay to only attention blocks and only MLP blocks within each transformer block. While our framework does work well and gives similar results in these settings, we do observe a performance drop using the same $\alpha_{\rm min}$. This can be explained by the reduction in the number of weighted connections. By halving the number of weighted residuals, an increase in the decay rate is required to match the quality of learned representations. Hence, our framework is actually robust against particular design choices of which connections are decayed but relies more on the total rate of information decay throughout the encoder.
We also explore a different parameterization of the residual connection. A few existing works leverage the benefits of re-scaling the residual connections in the context of image generation tasks [62, 63, 64, 65], but all of them scale both branches of the residual block jointly. Following this design, we try applying a constant factor $\alpha=\sqrt{0.5}$ to both ${\bm{x}}$ and $f_{{\bm{\theta}}_{l}}({\bm{x}})$, as seen in the implementation of Karras et al. [65]. In Table LABEL:tab:decay_schema, we observe that including such constant re-scaling to both branches decreases the linear probing score of the model. We also try applying a constant scaling factor to only $f_{{\bm{\theta}}_{l}}({\bm{x}})$, a case studied by [27]. Decaying $f_{{\bm{\theta}}_{l}}$ term does not alter the results much since the model can easily learn a scaling factor within $f_{{\bm{\theta}}_{l}}$ to counterbalance the added decaying force. These two ablations confirm that the key for abstract representation learning is to control the weights of the identity shortcuts, leveraging the structural bias reminiscent of feed-forward neural networks.
@@ 5 Discussion on feature rank
In this section, we try to answer a key question: How and why do residual connections impact the abstraction level of the deeper layers in a neural network? We delve deeper into how our design reinforces the low-rank bias of neural networks and try to connect our method to ideas in existing works [28]. To this end, we visualize the training dynamics of our method and analysis the feature rank of our approach to provide a holistic analysis.
Low-rank simplicity bias. Huh et al. [28] investigate the low-rank simplicity bias in deeper feed-forward neural networks, which drives neural networks to find low-rank solutions. At the same time, they make an empirical observation that deeper residual networks do not show a similar rank contracting behavior.
Effective rank. For analysis purpose, Huh et al. [28] quantify the rank of the learned representation using the effective rank, which is a continuous measure. For a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, the effective rank $\rho({\bm{A}})$ is defined as the Shannon entropy of the normalized singular values [66]:
@@
where $\bar{\sigma}_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$ denotes the $i^{\text{th}}$ normalized singular value. Intuitively, this measure is small when a few singular values dominate and large when singular values are evenly spread, hence giving a good continuous approximation for matrix rank.
In the following subsections, to compute the effective rank, we use the singular values from the covariance matrix ${\bm{A}}_{\bm{\theta}}$ of the last-layer features, where ${\bm{A}}_{\bm{\theta}}(i,j)$ denotes the covariance of the learned class tokens for the $i^{\text{th}}$ and $j^{\text{th}}$ samples.
@@
Figure 4: We visualize (a) the training dynamics of effective rank for various $\alpha_{\rm{min}}$ and (b) the linear probing accuracy for various $\alpha_{\rm min}$ for unsupervised Masked Autoencoders (MAE) trained on ImageNet-100. These features are extracted from the bottleneck layer of the MAE encoder. In (a) we observe that $\alpha_{\rm min}=0.6$ and $\alpha_{\rm min}=0.5$ produce features with the lowest effective rank. In (b) we observe that accuracy improves with lower $\alpha_{\rm min}$.
Inspired by their analysis, we conjecture that the improvement to feature learning capability of our method can mainly be attributed to the decayed identity shortcuts promoting low-rank features at the bottleneck. We measure the training dynamics of the models presented in Table 2 (MAEs trained on ImageNet-100) in terms of accuracy and the effective rank.
In Figure 4(a) and 4(b), we present the training dynamics of our model, illustrating effective rank and linear probing accuracy across various values of $\alpha_{\text{min}}$. During training, models with lower $\alpha_{\text{min}}$ tend to exhibit higher accuracy and lower effective rank, which supports the conjecture that low-rank features contribute to better learned representations in the MAE model. Additionally, we observe that these patterns persist throughout the training process, suggesting that establishing a low-rank bottleneck from the outset benefits model training right from the beginning.
Overall, we have uncovered evidence supporting our conjecture that an appropriate low-rank bias helps abstract representation learning and shortcut decay in residual connections promotes this bias, enhancing the abstract feature learning performance.
Compatibility with contrastive learning frameworks. Though seeing the big improvement of applying our decayed identity shortcuts in masked autoencoders, we note that our approach does not easily extend to contrastive learning frameworks, where the low rank inductive bias conflicts with the training objectives. For example, the training objective in MoCov3 [58] include an universal repulsive term in the denominator, which is supposed to increase the feature rank. Furthermore, contrastive learning also imposes specific data distribution requirements, such as both image crops needing to include the same object, which is impractical for in-the-wild images. In contrast, masked autoencoders do not have this requirement, offering greater potential for scalability and generalization. Our work focuses on further improving the learned representations of masked autoencoders, and enhancing contrastive learning is beyond the scope of our discussion.

### v3

@@ 3.2 Implementation Strategy
Skip connections for autoencoders. Since our method progressively decays the residual connections over network depth, it encourages the most abstract features to be learned by later layer. However, learning a highly abstract bottleneck is detrimental to the training objectives that aim for pixel-wise reconstruction, as they necessitate the preservation of information across all feature levels. To address this, we incorporate standard skip connections between the encoder and decoder, enabling the encoder to directly pass information from shallow layers to the decoder while learning increasingly abstract representations in the deeper encoder layers.
@@
Stabilizing training with residual zero initialization. The model exhibits rapid feature norm growth at the beginning of training for $\alpha_{\rm min}\leq 0.7$. We suspect that the model learns to amplify the output feature norm of $f_{{\bm{\theta}}_{l}}({\bm{x}})$ to counteract the significant decay applied to the residual connection. This growth leads to training instability and negatively impacts training convergence. To address this issue, we follow the implementation of previous works (Ho et al., 2020) and initialize the weights of the final output layer in each $f_{{\bm{\theta}}_{l}}$ to zero instead of using the original Xavier uniform initialization (Glorot & Bengio, 2010). This approach significantly enhances training stability by limiting the rate of feature norm growth and enables us to explore training with even lower values of $\alpha_{\rm min}$.
4 Experiments on Masked Autoencoder (MAE)
@@
For masked autoencoders (MAEs) (He et al., 2022), we replace the residual connections in the encoder’s MLP and attention blocks with decayed identity shortcuts. The MAE operates by accepting images with a random subset of pixels masked out and learning to recover the discarded pixels. As Section 3.2 describes, we add skip connections between the encoder and decoder to facilitate learning abstract representations at the bottleneck. Since the original MAE has twice the number of encoder layers as decoder layers, we inject output from every other encoder layer into the corresponding decoder layer. To match spatial dimensions, injected encoder features are combined with learnable masked tokens before channel-wise concatenation. The implementation details for the training and evaluation are shown in Section A. He et al. (2022) show the desired representations appear at the end of encoder; we therefore apply our decaying schema only to the encoder.
4.1 Representation Learning on ImageNet-1k
We pre-train MAE on the ImageNet-1K train split (Deng et al., 2009) and follow recent works to evaluate the learned representations using linear probing and end-to-end finetuning, where we added a single linear head over learned representations to predict the image categories
For hyperparameters of pretraining MAE, including the learning rate schedule, total training epochs, and mask ratio, we follow the best settings found in the original paper. Please see the appendix for detailed experimental setups.
Method Objective Augmentation FT LP
MoCo-v3(Chen et al., 2021) Contrastive Loss Full 83.2 76.7
DINO(Caron et al., 2021) Contrastive Loss Full 83.3 78.2
Con MIM(Yi et al., 2023) Contrastive Loss Affine + Mask 83.7 39.3
ADDP(VIT-L) (Tian et al., 2024) Feature Loss (VQ) Affine + Mask 85.9 23.8
Latent MIM(Wei et al., 2024) Feature Loss Affine+ Mask 83.0 72.0
Data2Vec(Baevski et al., 2022) Feature Loss Full+Mask 84.2 68.0
CAE(Chen et al., 2024) Recon. + Feature Loss Affine+Mask 83.8 70.4
I-JEPA(Assran et al., 2023) Feature Loss Affine+Mask - 72.9
MAE(He et al., 2022) Recon. Loss Affine+Mask 83.6 67.8
Ours ($\alpha_{\rm min}$ = 0.6) Recon. Loss Affine+Mask 82.9 72.7
@@ 4.1 Representation Learning on ImageNet-1k
Table 1: Accuracy of linear classifier based on pre-trained learned representations on the ImageNet-1K dataset. We evaluate our learned representation using the standard evaluation protocol: linear probing (LP) and fine-tuning (FT). With our simple modification, we substantially improve the MAE ViT-B/16 baseline for linear probing by encouraging the model to learn an increased abstract features over depth. We achieve competitive performance compared to I-JEPA which uses an explicit feature loss.
@@
We report the results in Table 1, where we show the linear probing performance of various self-supervised methods, which we categorize by their objectives, and data augmentation processes. In the top half of the table, we present methods that employ a contrastive loss. Although these methods produce the best probing accuracies, their success depends on a carefully designed data augmentation process, which may need to be tuned for each different data distribution. In the bottom half, we show several methods based on generative architecture, including ours, which do not rely on contrastive objectives. With the exception of Data2Vec, these methods only use a standard random affine data transformation (with random masking), which need not be distribution-reliant. Among these methods, MAE only uses a pixel-wise loss, I-JEPA , Latent MIM, and CAE use a latent feature alignment loss, and CAE uses both. Our method simply extends MAE by constructing an implicit feature bottleneck and shows significant improvements over the MAE baselines (72.7% vs. 67.3%), outperforming Data2Vec, Latent MIM and CAE and giving a probing accuracy competitive with I-JEPA, without needing to use explicit feature alignment.
End-to-end fine-tuning, unlike linear probing which only trains a single linear layer, updates the entire network for image classification. Since the features can shift significantly from their pre-training state during end-to-end updating, we argue that this may not accurately reflect the quality of the learned representations. For example, DINO demonstrates superior performance in various downstream vision tasks compared to MAE, but its fine-tuning performance is worse than MAE. Similarly, ConMIM and ADDP exhibit poor linear probing performance, suggesting lower-quality representations, yet their fine-tuning performance surpasses that of contrastive learning methods. Nevertheless, we still provide the fine-tuning results for reference.
(a) MAE
(b) Ours
@@
Figure 3: Comparison of t-SNE visualization for models trained with (a) standard residual connection and (b) our method with $\alpha_{\rm min}=0.6$. We visualize the learned embedding using t-SNE on 10 randomly selected categories of ImageNet-100 and we color points within the same category using the same color. Model with standard residual connections (a) have collapsed together while our method (b) forms well structured clusters.
(c) Input Image
(d) MAE (4.1 mIoU)
(e) Ours (10.4 mIoU)
Figure 4: Visualize learned representations using Zhang et al. (2024). We project the learned representations onto a 3-channel feature map, visualized as RGB images. Our method learns more abstract and semantically consistent representations compared to the baseline MAE. This visual comparison is further supported by benchmarking on unsupervised semantic segmentation tasks, where our approach achieves better results (10.4 mIoU) compared to the baseline MAE (4.1 mIoU).
4.2 Embdding analysis.
@@
We qualitatively evaluate the feature learning in Figure 3(a) by visualizing the t-SNE embeddings of the learned features for a random subset of test images. The embedding produced from our features displays much clearer separation than baseline, which has struggles to differentiate the categories.
To provide another qualitatively evlauation of our learned representations, we adopt the pixel-wise embedding approaches proposed by Zhang et al. (2024) to group the representations from the last layer of the encoder into a lower dimensional space. We use their default hyperparameters to cluster representations across images of COCO validation set. We visualize in Figure 4 the top 3 eigenvectors for both our approaches with $\alpha_{\min}=0.7$ and baseline. From the visualization, ours learns abstract representation and the object from the same categories have similar color, indicating a global consistent semantic grouping. The baseline MAE, on the other hand, doesn’t show clearly global semantic patterns overall and the low-level representations, e.g., edges, are also high-lightened, indicating a representations from multiple semantic hierarchies are entangled.
We further evaluate the clustering quantitatively, following the postprocessing protocol  (Zhang et al., 2024) to produced unsupervised semantic segmentation and report the results as the mean intersection of union (mIoU). Ours (10.41 mIoU) achieve 6.31 mIoU improvement over baseline (4.10 mIoU), which support the qualitatively comparison.
4.3 Ablation Studies on ImageNet-100
@@
Table 2: Linear Probing Accuracy on ImageNet-100 for our method varying $\alpha_{\rm min}$ and architecture size. We conduct ablation studies and demonstrate that linear probing performance for both architectures increases as $\alpha_{\rm min}$ decreases until around 0.5-0.6. While the larger ViT-B/16 architecture achieves the highest accuracy of 83.6, it is noteworthy that the smaller ViT-S/16, when utilizing our method, outperforms the baseline setting (standard Residual Connection at $\alpha_{\rm min}=1$) of ViT-B/16.
@@
Architecture size. In Table 2, we also train both the ViT-B/16 and the smaller ViT-S/16 backbone using varying $\alpha_{\rm min}$. Our framework is especially effective on the smaller architecture, increasing linear probing performance by 9.4%, compared to the baseline setting (standard Residual Connection at $\alpha_{\rm min}=1$). This is consistent with the observation that larger models generalize better (Huh et al., 2021) than smaller models. Our method is able to significantly improve the smaller ViT-S/16 and slightly close the gap between the differently-sized architectures.
Skip connections. Another critical design choice in our network is to include skip connections that are not in the original MAE. As discussed in Section 3.2, if the MAE does not use skip connections, the bottleneck layer must preserve all information to reconstruct the input image accurately. This is opposed to learn abstract representations at bottleneck. These contrary effects significantly degrade the representation learned by the model, leading to a 22.1% drop in the linear probing score, as we report in Table LABEL:tab:skip_connections.
Different decay schema. We also explore decay schema, with results summarized in Table LABEL:tab:decay_schema: (1) Scaling both branches of the residual blocks simultaneously by applying a constant factor, $\alpha=\sqrt{0.5}$, to both ${\bm{x}}$ and $f_{{\bm{\theta}}_{l}}({\bm{x}})$. (2) Scaling only $f_{{\bm{\theta}}_{l}}$ using the same constant factor, $\alpha=\sqrt{0.5}$. (3) Applying our proposed schema exclusively to either the attention or MLP branch.
Among these, (2) shows no significant improvement over the baseline, while (1) yields some improvement but still underperforms compared to our approach. By analyzing (1) and (2), we demonstrate that the representation gains are due to down-weighting the skip connection branch. Notably, recent diffusion models (Karras et al., 2018; Song et al., 2021b; Karras et al., 2020; 2022) have employed (1) in their designs. However, applying decay only to the MLP or attention branch reduces the overall decaying effect across the network, resulting in lower performance compared to our schema, which achieves the best performance among the tested designs.
@@
5 Experiments on Diffusion Models
Diffusion models. We use U-ViT (Bao et al., 2023), a ViT-based diffusion model with skip connections between the encoder and decoder, as the baseline for our diffusion model experiments. Recent studies (Yang & Wang, 2023; Baranchuk et al., 2022) suggest that diffusion models learn the best semantic representations near the decoder’s latter stages. Therefore, we apply our proposed decay mechanism up to the end of the decoder. While this design might be suboptimal, as the smallest decay factor may not align with the layers holding the best semantic representations, we demonstrate in practice that this simple approach effectively enhances both the learned representations and the quality of generated outputs.
Experimental details. We utilize the default scheduler and sampler from U-ViT (Bao et al., 2023), replacing only the residual connections with our proposed decayed shortcut connections. We train unconditional diffusion models on CIFAR-100 and ImageNet-100 without using image class labels. Additionally, we train a class-conditional diffusion model on ImageNet-100 to validate our design across different tasks. For ImageNet-100, instead of training directly on pixels, we adopt a latent diffusion (Rombach et al., 2022) approach by running the model in the latent space of a pretrained VAE, which reduces input resolutions from 256x256x3 to 32x32x4. We use U-ViT-Mid for ImageNet-100 and U-ViT-small for CIFAR-100. For model and training details, please refer to  Bao et al. (2023).
We evaluate the learned representations with linear probing and we train a linear classifier over the frozen representations. We report the results as the best configurations, including the choices of layer index and noise level, that yields the best performance
Results. Our results are presented in Table 4, where we demonstrate that replacing residual connections with our proposed decayed identity shortcuts consistently enhances representation quality and image generation across both datasets and tasks (conditional and unconditional generation). Notably, this improvement is achieved without introducing any additional learnable parameters. We provide the visualization of generated images in the appendix for qualitative comparisons (Figure 10).
Linear Probing (Acc.$\uparrow$ Generation quality (FID)$\downarrow$
Dataset $\alpha_{\rm min}$ 1.0 0.8 0.7 0.6 1.0 0.8 0.7 0.6
CIFAR-100 (Uncon.) 62.47 63.58 66.86 64.63 14.34 11.65 8.99 11.71
ImageNet-100 (Uncond.) 72.8 74.5 76.1 75.8 44.40 40.96 41.17 43.51
ImageNet-100 (Class Cond.) - - - - 6.93 5.75 5.11 4.98
Table 4: Performance using diffusion models. In diffusion models, we demonstrate that our proposed decayed identity shortcut (with $\alpha_{\rm min}<1.0$) enhances probing accuracy and improves generation quality across various datasets and configurations in both classes conditional and unconditional setups.
@@ 6 Discussion on feature rank
In this section, we try to answer a key question: How and why do residual connections impact the abstraction level of the deeper layers in a neural network? We delve deeper into how our design reinforces the low-rank bias of neural networks and try to connect our method to ideas in existing works (Huh et al., 2021). To this end, we visualize the training dynamics of our method and analysis the feature rank of our approach to provide a holistic analysis.
Low-rank simplicity bias. Huh et al. (2021) investigate the low-rank simplicity bias in deeper feed-forward neural networks, which drives neural networks to find low-rank solutions. At the same time, they make an empirical observation that deeper residual networks do not show a similar rank contracting behavior.
Effective rank. For analysis purpose, Huh et al. (2021) quantify the rank of the learned representation using the effective rank, which is a continuous measure. For a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, the effective rank $\rho({\bm{A}})$ is defined as the Shannon entropy of the normalized singular values (Roy & Vetterli, 2007):
@@
where $\bar{\sigma}_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$ denotes the $i^{\text{th}}$ normalized singular value. Intuitively, this measure is small when a few singular values dominate and large when singular values are evenly spread, hence giving a good continuous approximation for matrix rank. In the following subsections, to compute the effective rank, we use the singular values from the covariance matrix ${\bm{A}}_{\bm{\theta}}$ of the last-layer features, where ${\bm{A}}_{\bm{\theta}}(i,j)$ denotes the covariance of the learned class tokens for the $i^{\text{th}}$ and $j^{\text{th}}$ samples.
@@
[…]
T
[…]
coefficient
$r_{\textrm{rank, accuracy}}$
200 -0.51
300 -0.56
400 -0.57
500 -0.58
600 -0.52
700 -0.44
800 -0.42
(c) Pearson correlation coefficient between effective rank and probing accuracy, across training epochs.
Figure 6: For MAE pretrained on ImageNet-100, we present visualizations of (a) the training dynamics of the effective rank for different values of $\alpha_{\rm{min}}$, (b) the linear probing accuracy for various $\alpha_{\rm{min}}$ , and (c) the Pearson correlation coefficient between feature rank and probing accuracy, demonstrating that a lower effective feature rank is associated with better performance.
Inspired by Huh et al. (2021), we conjecture that the improvement to feature learning capability of our method can mainly be attributed to the decayed identity shortcuts promoting low-rank features at the bottleneck. We measure the training dynamics of the models presented in Table 2 (MAEs trained on ImageNet-100) in terms of accuracy and the effective rank. In Figure 6(c), we quantify the correlation between effective feature rank and probing accuracy using Pearson correlation coefficient to validate the hypothesis.
In Figures 6(a) and 6(b), we present the training dynamics of our model, highlighting effective rank and linear probing accuracy for different values of $\alpha_{\text{min}}$. During the early epochs, models with lower $\alpha_{\text{min}}$ tend to exhibit both lower effective rank and higher probing accuracy, supporting our hypothesis. As training progresses, the correlations between $\alpha_{\text{min}}$ and effective rank become less precise. We suspect this is due to the training dynamics, where models with lower $\alpha_{\text{min}}$ experience faster growth in feature rank, reflecting the model’s effort to compensate for the decay factors. Despite this, we can still conclude that lower $\alpha_{\min}=[0.5-0.6]$ results in lower feature rank and better probing accuracy compared to higher $\alpha_{\min}=[0.7-1.0]$.